OpenStudy (anonymous):

I have a question with cubed root and inside the radical symbol is 64(x=3)^3 I tried to work it out just need a little help to verify.

OpenStudy (anonymous):

this is what the problem looks like

OpenStudy (anonymous):

my answer was 4x+12

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{64} \times \sqrt[3]{(x+3)^3}\] Hence, 4(x+3)

OpenStudy (saifoo.khan):

So your answer is correct.

OpenStudy (anonymous):

yay! =)

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

Np